Determine the gas flow rate (Nm3 /m2 -min) in upper zone of a blast furnace for the given data: permeability = 2.2x10-6 m4 /N-sec, suction (AP) = 1000 mm of water, bed depth (H) = 50 cm. 


1
Expert's answer
2021-02-05T04:01:46-0500

Flow rate

= 2.2 × 10 -6 × 1000 / 10 × 50

= 2.2 × 500 × 10-6 .

= 1100 × 10 -6 ..